OpenStudy (anonymous):

solve and check help plezz 2x+5<8 the answer is x<3/2 how do i check this??

OpenStudy (mitu12):

mathway.com

OpenStudy (mitu12):

your answer is correct

OpenStudy (anonymous):

yes i kno but i have to provide the check thing

OpenStudy (anonymous):

can u help?

OpenStudy (texaschic101):

2x + 5 < 8 2x < 8 - 5 2x < 3 x < 3/2 now we know that x is less then 3/2...so pick a number less then 3/2 (1 1/2) such as 1 1/4, which is 5/4 now sub 5/4 in for x in the inequality to see if it is true 2x + 5 < 8 2(5/4) + 5 < 8 5/2 + 5 < 8 5/2 + 10/2 < 8 15/2 < 8 7 1/2 < 8 (correct)

OpenStudy (texaschic101):

if you picked a number greater then 8, then the inequality would be false.

OpenStudy (texaschic101):

oops...picked a number greater then 3/2, it would be incorrect

OpenStudy (anonymous):

so instead of 7 1/2<8 it should be 3/2?

OpenStudy (texaschic101):

the 7 1/2 < 8 was my way of checking that the inequality was true when subbing in 5/4, a number less then 3/2.

OpenStudy (texaschic101):

you should be able to sub in any number less then 3/2 for the inequality to be correct.

OpenStudy (anonymous):

ohh i see. i have another question how wud i check x^2-3x+5=x(x-6)+20 my answer is x=5

OpenStudy (texaschic101):

you would check it by subbing in 5 for x x^2 - 3x + 5 = x(x - 6) + 20 5^2 - 3(5) + 5 = 5(5 - 6) + 20 25 - 15 + 5 = 5(-1) + 20 30 - 15 = -5 + 20 15 = 15 (correct) they equal and you checked it by subbing in 5 for x

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

okay i have 2 more. god ur a life savor. 3x+4>10 my answer is x>2